### v5

@@ 4.3 Overall flowchart of this LFCO model
Tables 1 and 2 briefly presents the categorical meanings of the graphical elements (nodes and arrows) as ‘CLASSes’ in Figs. 6, 10 and 11. The physical/mathematical significance of the specific instantiated ‘OBJECTs’ within each ‘CLASS’ can be found in Supplementary Notes 2-5.
Starting from the atomic inputs, we built the comprehensive workflow diagram in Fig. 11 from the bottom up, referencing three submodules from Fig. 10a,c. All elements in the flowcharts (Figs. 6, 10 and 11) of this article are accompanied by step-by-step derivations and detailed explanations provided in the Supplementary Notes.
@@ Acknowledgements
We thank Sir M.V. Berry for the insights shared through both our online and offline exchanges. Special thanks go to “Mr. Belly” for honoring NJU-CEAS with his e(n)lightening public lecture-seminar trilogy, completing the trilogy of trilogies. We are also grateful to the editor and reviewers for their constructive and tolerant criticism of the initial manuscript, helping evolve its progression from a pure mathematical exploration of CO to a trinity of CO, NO, and FO. The first author extends deep thanks to all who shared a “light-matter interaction” with this monk during his three-year immersion in crafting the trilogy — without them, it would not have come into being.
@@
Funding This work was supported by the Natural Science Foundation of Jiangsu Province (BK20240005) and the Fundamental Research Funds for the Central Universities (021314380268). 1 person, 2 funds, 3 years111Curiously enough, this work came into being precisely on the occasion of NJU’s 123rd anniversary.? Oliver wants more. Oliver asserts that it may lead to some (Ig) Nobel (level) Prizes, should his coherent vision can be markedly propelled by substantial external funding. In addition, Oliver keeps a list of well-examined individuals and projects, which parallel this trilogy and all it has rigorously assessed — each still at a crescent phase.

### v6

@@ 4.3 Overall flowchart of this LFCO model
Tables˜1 and 2 briefly presents the categorical meanings of the graphical elements (nodes and arrows) as ‘CLASSes’ in Figs.˜6, 10 and 11. The physical/mathematical significance of the specific instantiated ‘OBJECTs’ within each ‘CLASS’ can be found in Supplementary Notes 2-5.
Starting from the atomic inputs, we built the comprehensive workflow diagram in Fig.˜11 from the bottom up, referencing three submodules from Fig.˜10a,c. All elements in the flowcharts (Figs.˜6, 10 and 11) of this article are accompanied by step-by-step derivations and detailed explanations provided in the Supplementary Notes.
@@ Acknowledgements
We thank Sir M.V. Berry for the insights shared through both our online and offline exchanges.
@@
Funding This work was supported by the Natural Science Foundation of Jiangsu Province (BK20240005) and the Fundamental Research Funds for the Central Universities (021314380268).